# Open Materials Generation with Stochastic Interpolants

### v1

@@ C.1 Match rate and RMSE
We include here a discussion on the tradeoff between match rate and RMSE, as we noticed this most strongly influences the perov-5 dataset. We show in Fig. 7 how different interpolants for the atomic coordinates (trigonometric vs. linear) learn to match structures differently. For the linear case (which has a depressed match rate) the change in matching tolerance makes little difference; for the trigonometric interpolant it makes a far more significant difference and leads to a much higher match rate, suggesting that the trigonometric interpolant learns structures more reliably but less accurately.
@@
Figure 7: We show here the effect of making matching more difficult by decreasing the length tolerance used by pymatgen’s StructureMatcher. We plot the density of the normalized RMSE distributions from CSP models trained on the perov-5 dataset. We note that the curves for all generated structures and only valid generated structures overlap significantly.
C.2 Validity metrics
@@
We noted that for the compositional validity metric, the version of the SMACT software [43] changed the computed values by several percent. This impacts both the match rate (when filtered by valid structures) as well as the reported DNG compositional validity output. As such, all values for all models were recomputed with the most up-to-date version (3.0) of the SMACT software.
C.3 Calculation of S.U.N. rates
@@
Evaluation of DNG structures was performed using scripts provided by the developers of MatterGen [53]. A total of 10,000 structures were generated from each of OMG, DiffCSP, and FlowMM. These structures were then filtered to remove any that contained elements not supported by the MatterSim potential (version MatterSim-v1.0.0-1M) or the reference convex hull. These included heavy elements with atomic numbers $>$89, radioactive elements, and the noble gases (specifically: ‘Ac’, ‘U’, ‘Th’, ‘Ne’, ‘Tc’, ‘Kr’, ‘Pu’, ‘Np’, ‘Xe’, ‘Pm’, ‘He’, ‘Pa’). Stability and novelty were computed with respect to the default dataset provided by MatterGen which contains 845,997 structures from the MP-20 [55, 29] and Alexandria [58, 57] datasets. This provides a more challenging reference for computing novelty as each model was trained only on the $\sim$ 27,000 structures from the MP-20 training set.
C.4 Structural and stability analysis of generated structures
@@
In Fig. 8, we show the distribution of computed energies above the convex hull across various OMG models, showing best stability of generated structures for linear, encoder-decoder, trigonometric, and score-based diffusion (SBD) positional interpolants.
Figure 8: Histogram of the computed energies above the convex hull for structures generated by four OMG DNG models trained using joint evolution for cell, coordinates, and species. We show that the linear and encoder-decoder positional interpolants are effective at generating more structures close to the convex hull. See Appendix C.3 for calculation details.
@@
By evaluating the distribution of $N$-ary structures (Fig. 9), the distribution of average coordination numbers (both by structure in Fig. 10 and by species in Fig. 11), as well as distribution of crystal systems (Fig. 12) which are related to a structure’s Bravais lattice, we provide qualitative analysis for model performance. Space groups (and thus crystal system) were determined using the spglib software [66]. Generally, OMG and DiffCSP showed the best performance in matching the distribution of average coordination number for each structure, particularly for high-coordinated structures. DiffCSP performed most poorly among the four models on the $N$-ary structural similarity. The average coordination number for species were best-matched by the two OMG models, and of these two the linear interpolant generated the distribution of crystal systems best.
Figure 9: Qualitative performance of the distribution of $N$-ary crystals for (a) Non-OMG models and (b) OMG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
Figure 10: Qualitative performance of the distribution of average coordination number by structure for (a) Non-OMG models and (b) OMG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
@@
Figure 11: Qualitative performance of the distribution of average coordination number by species (listed left to right in order of atomic number) for (a) Non-OMG models and (b) OMG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
@@
Figure 12: Qualitative performance of the distribution of crystal system by structure for (a) Non-OMG models and (b) OMG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.

### v2

@@ D.1 Match Rate and RMSE
The tradeoff between match rate and RMSE most strongly influences the perov-5 dataset. We show in Fig. 7 how different positional interpolants for the atomic coordinates (trigonometric vs. linear with ODE sampling schemes) learn to generate matched structures differently. For the linear case, the change in matching tolerance (via the ltol parameter of Pymatgen’s StructureMatcher) makes little difference. For the trigonometric interpolant, it makes a far more significant difference and leads to a much higher match rate, suggesting that the trigonometric interpolant learns structures more reliably but less accurately.
@@
Figure 7: We show here the effect of making matching more difficult by decreasing the length tolerance used by Pymatgen’s StructureMatcher. We plot the density of the normalized RMSE distributions from CSP models trained on the perov-5 dataset (Linear and Trigonometric positional interpolants with ODE sampling schemes). We note that the curves for all generated structures and only valid generated structures overlap significantly.
D.2 Validity Metrics
@@
The structural validity of generated structures is defined according to the bond lengths present in the structure—all lengths must be $>$0.5 Å to be considered valid. The compositional validity is defined according to the SMACT software package [45]. We note that the default oxidation states have been updated with the release of SMACT version 3.0 which changed the DNG compositional validity rates by several percent. This also impacts the CSP match rate when filtered by valid structures. As such, all values for all models were recomputed with the most up-to-date version (3.0) of the SMACT software.
D.3 Coverage and Property Statistics
As in [29], we evaluate coverage recall and precision (reported as rates) by measuring the percentage of crystals in the test set and in the generated samples that match each other within a defined fingerprint distance threshold. For structural matching, we use CrystalNN, and for compositional matching, we use Magpie fingerprints. Additionally, we report Wasserstein distances between property distributions of the generated and reference datasets. The considered properties are the mass density ($\rho$), the number of unique elements ($N$ary), and the average coordination number for each element in the unit cell ($\langle CN\rangle$).
D.4 Calculation of S.U.N. Rates
Table 12: Stability (defined as $\leq 0.1$ eV/atom above hull), uniqueness, and novelty results from de novo generation on the MP-20 dataset computed for the same models as in Tab. 2. All evaluations are performed with the MatterGen code base [53] with respect to the included reference Alex-MP-20 dataset. The average RMSD is between the generated and the relaxed structures, and the average energy above hull is reported in units of eV/atom. All structures were relaxed with the MatterSim potential. Note that the results in Tab. 3 relied on a subsequent DFT relaxation.
Method $\langle E\rangle/N$ ($\downarrow$) RMSD Novelty Stability S.U.N.
above hull (Å, $\downarrow$) Rate (%, $\uparrow$) Rate (%, $\uparrow$) Rate (%, $\uparrow$)
DiffCSP 0.1984 0.367 72.73 43.04 19.00
FlowMM 0.2509 0.651 72.76 37.47 13.86
MatterGen-MP 0.1724 0.142 72.17 47.07 22.66
OMatG
Linear 0.1823 0.615 72.00 45.00 22.07
Trig 0.1857 0.657 65.35 51.40 19.96
Enc-Dec 0.1699 0.390 54.97 58.56 17.59
SBD 0.2189 0.763 75.80 42.60 22.10
CFP + CSP 0.2340 0.488 75.85 42.21 20.50
@@ D.3 Coverage and Property Statistics
Evaluation of DNG structures was performed using scripts provided by the developers of MatterGen [53]. A total of 10,000 structures were generated from each of OMatG, DiffCSP, FlowMM, and MatterGen-MP. These structures were then filtered to remove any that contained elements not supported by the MatterSim potential (version MatterSim-v1.0.0-1M) [54] or the reference convex hull. These included heavy elements with atomic numbers $>$89, radioactive elements, and the noble gases (specifically: ‘Ac’, ‘U’, ‘Th’, ‘Ne’, ‘Tc’, ‘Kr’, ‘Pu’, ‘Np’, ‘Xe’, ‘Pm’, ‘He’, ‘Pa’). [^7] These elements were not removed from any datasets during training. Stability and novelty were computed with respect to the default dataset provided by MatterGen which contains 845 997845997845\,997845 997 structures from the MP-20 [56, 29] and Alexandria [59, 58] datasets. This provides a more challenging reference for computing novelty as each model was trained only on the 27 000similar-toabsent27000$\sim$\$27\,000\$∼ 27 000 structures from the MP-20 training set.
The MatterSim potential was utilized for the first structural relaxation, requiring far less compute resources compared to full DFT. Results derived from the MatterSim-relaxed structures are shown in Tab. 12. Following the MatterSim relaxtions, 1000 structures were relaxed with DFT. All DFT relaxations utilized MPGGADoubleRelaxStatic flows from the Atomate2 [71] package to produce MP20-compatible data.
Comparing the results from Tab. 12 to Tab. 3, we find that overall there is reasonable agreement between the metrics computed at the machine learning potential and DFT level. Relative performance ordering between methods remains fairly consistent, allowing for qualitative trends to be made at the much cheaper ML potential level. Nevertheless, for a full quantitative understanding, DFT is essential.
D.5 Stability and Structural Analysis of Generated Structures
@@
In Fig. 8, we show the distribution of computed energies above the convex hull across various OMatG models, showing best stability of generated structures for linear, encoder-decoder, trigonometric, and VP SBD positional interpolants.
Figure 8: Histogram of the computed energies above the convex hull for structures generated by the five OMatG DNG models highlighted in the main text (see Table 3). We show that all positional interpolants are effective at generating structures close to the convex hull, with the VP SBD interpolant and CFP + CSP method performing slightly worse than other interpolants.
@@
By evaluating the distribution of $N$ary structures (Fig. 9), the distribution of average coordination numbers (both by structure in Fig. 10 and by species in Fig. 11), as well as distribution of crystal systems (Fig. 12) which are related to a structure’s Bravais lattice, we provide qualitative analysis for model performance. Space groups (and thus crystal systems) were determined using the spglib software [72] and choosing the most common space group identification with exponentially (geometrically) decreasing tolerance. We find that OMatG models have superior performance on the $N$ary metric across the board, with DiffCSP performing the most poorly. Specific OMatG models (with positional Encoder-Decoder interpolant and CFP+CSP with Linear interpolant) and DiffCSP showed the best performance in matching the distribution of average coordination number for each structure, particularly for high-coordinated structures. The average coordination number for species were best-matched by OMatG models across the board (with the exception of the OMatG-VPSBD model which tended to overpredict the coordination environments), and broadly underpredicted by DiffCSP, FlowMM, and MatterGen-MP. Finally, the best matching results for distribution across crystal systems was for the OMatG CFP+CSP model, with all other interpolants (and DiffCSP and FlowMM) showing a propensity for generating low-symmetry (triclinic and monoclinic) crystal structures. Overall, we find that OMatG models closely reproduce the elemental and structural diversity present in the data.
@@
Figure 9: Qualitative performance of the distribution of $N$ary crystals for (a) Non-OMatG models and (b) OMatG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset. Atomic elements are listed in increasing atomic number from left to right.
@@
Figure 10: Qualitative performance of the distribution of average coordination number by structure for (a) Non-OMatG models and (b) OMatG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
Figure 11: Qualitative performance of the distribution of average coordination number by species (listed left to right in order of atomic number) for (a) Non-OMatG models and (b) OMatG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
Figure 12: Qualitative performance of the distribution of crystal system by structure for (a) Non-OMatG models and (b) OMatG models across structural benchmarks computed on generated structures and test set structures from the MP-20 dataset.
Appendix E Large Language Models as Base Distributions
FlowLLM [61] combines large language models (LLMs) with the conditional flow-matching framework FlowMM [32] to design novel crystalline materials in the DNG task. A fine-tuned LLM serves as the base distribution and samples initial structures; FlowMM then refines the fractional coordinates and lattice parameters as in the CSP task. This idea can be similarly applied to OMatG, which allows the use of LLMs within the general SI framework for materials generation.
We extend OMatG to OMatG-LLM by allowing for LLM-generated structures as the initial structures. We evaluate both FlowLLM and OMatG-LLM on the LLM dataset released by FlowLLM. Specifically, we use the training (containing 40 0004000040\,00040 000 structures) and validation sets ($6000$ structures) from https://github.com/facebookresearch/flowmm and the LLM-generated initial structures (10 0001000010\,00010 000 structures) from https://github.com/facebookresearch/crystal-text-llm as the test set. These initial structures are generated by a fine-tuned Llama-70B model [73]. As shown in Tab. 13, OMatG-LLM’s linear and trigonometric interpolants outperform FlowLLM in almost all DNG metrics. Since the Wasserstein distance with respect to the $N$arity distributions and the compositional validity only depend on the atomic species generated by the LLM, these two metrics are necessarily equal for FlowLLM and OMatG-LLM. Note that the original FlowLLM [61] is trained on 3M LLM-generated structures while our experiments are conducted on the 40K structures open-sourced by the authors. The performance of FlowLLM in our experiments thus differs from the scores reported in [61].
Table 13: FlowLLM’s and OMatG-LLM’s performance (with linear and the trigonometric interpolants) when using the same fine-tuned LLM [73] as the base distribution. The best performance for each metric is in bold.
Method Validity ($\uparrow$) Coverage ($\uparrow$) Property ($\downarrow$) S.U.N. Rate ($\uparrow$)
FlowLLM 96.27 86.40 83.55 97.98 96.55 0.9922 0.5427 0.5936 10.28
OMatG-LLM
Linear 97.86 86.40 84.85 99.16 98.40 0.9100 0.5427 0.8600 12.61
Trigonometric (ODE) 95.70 86.40 83.25 98.57 98.24 0.7410 0.5427 0.6165 11.14
Trigonometric (SDE) 97.78 86.40 84.72 97.41 99.12 3.6214 0.5427 0.4448 11.86
Appendix F Computational Costs
In Tabs 14 and 15, we present the computational costs for both the CSP and DNG tasks. We compare the cost of training and integrating OMatG on the MP-20 dataset and show low computational costs for OMatG’s ODE scheme for both training and inference. The SDE scheme is more expensive but competitive. For these experiments, we use an Nvidia RTX8000 GPU with a batch size of 512 and 1000 integration steps.
Table 14: Computational costs for DiffCSP, FlowMM, OMatG (ODE) and OMatG (SDE) models trained on the CSP task.
Task OMatG (ODE) FlowMM OMatG (SDE) DiffCSP
Training (s / epoch) $56.8\pm 0.75$ $70.35\pm 1.38$ $89.0\pm 1.41$ $21.89\pm 0.31$
Sampling (s / batch) $313.67\pm 9.29$ $424.125\pm 11.78$ $479.5\pm 13.5$ $338.11\pm 11.93$
Table 15: Computational costs for DiffCSP, FlowMM, OMatG (ODE) and OMatG (SDE) models trained on the DNG task.
Task OMatG (ODE) FlowMM OMatG (SDE) DiffCSP
Training (s / epoch) $75.26\pm 2.08$ $73.32\pm 0.47$ $102.65\pm 1.87$ $21.85\pm 0.36$
Sampling (s / batch) $473.14\pm 13.20$ $469.93\pm 6.12$ $617.2\pm 18.2$ $322.63\pm 10.28$